[ the side of a cube measures 6 units. find the surface of the cube ]

Found 2 solutions by TimothyLamb, Alan3354:
Answer by TimothyLamb(4379)

a = 6ss
s = 6 units
---
a = 6^3 = 216 sq.units
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by Alan3354(69443)

[ the side of a cube measures 6 units. find the surface of the cube ]
----------
A cube has 6 faces.
The area of each is s^2, where s = the length of the sides.